Tar 25 27 20 24 20 20 21 24 CO 18 16 16 16 16 16 14 17 Nicotine 1.5 1.7 1.1 1.6 1.1 1.0 1.2 1.4 Cigarette Tar and Carbon Monoxide Refer to the table of data given in Exercise 1 and use the amounts of tar and carbon monoxide. a. Construct a scatterplot using the horizontal scale to represent the amount of tar. What does the scatterplot suggest? b. Does the scatterplot show eight points corresponding to the eight pairs of sample data? If not, why not? c. Find the value of the linear correlation coefficient and determine whether there is sufficient evidence to support a claim of a linear correlation between amount of tar and amount of carbon monoxide. d. Letting y represent amount of carbon monoxide and letting x represent amount of tar, find the regression equation. e. The Raleigh brand king size cigarette is not included in the table, and it has 23 mg of tar. What is the best predicted amount of carbon monoxide? How does the predicted amount com- pare to the actual amount of 15 mg of carbon monoxide? Please show using Excel. Thank you
Answer #1

a) scttwr plot which is above

b) Scatter plot doesn't shows 8 pairs of sample data because for Tar vs CO, 2 pairs have same coordinates so it coinside and for Tar vs Nicotine , 1 pair have same coordinate. Points are present but due to coinsidation they shows only one point .

c) linear correlation coefficient is 0.4369 and there is evidence to support a claim of a linear correlation between amount of tar and amount of carbon monoxide.

d) linear regression equation - CO = 12.0361 + 0.1807 * Tar

e) Tar = 23 then CO = 12.0361 + 0.1807 * 23 = 16.1928

Actual value - predicted value = error , 15- 16.1928 = 1.1928 is the error . PL??